# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ B.3 Proof of Proposition 19
[…]
where the first convergence follows from the assumptions on $\log\widehat{F}(x,y\mid z_{1},z_{2})$ and $\log\widehat{G}(x,y\mid z_{1},z_{2})$. Note that $\widehat{\omega}_{n}$ is finite given Assumption A6 in Vuong (1989). We can characterize its behavior under the large sample limit as $\widehat{\omega}_{n}=O(1)$, since $\widehat{\omega}_{n}\rightarrow\omega^{*},\omega^{*}\in\mathbb{R}$. As $n\rightarrow\infty$, $\frac{1}{n}LR_{n}(\widehat{\theta}_{n},\widehat{\gamma}_{n})$ converges in probability to the true likelihood ratio statistic,
[…]
When $H_{f}$ is true, $c>0$; when $H_{g}$ is true, $c<0$. All together, the test statistic in Equation (26) can be written as $\sqrt{n}c\cdot O(1)$ and thus converges in probability to +$\infty$+ ∞ or $\infty$- ∞, respectively, under $H_{f}$ or $H_{g}$.
B.4 Proof of Theorem 2
@@
B.4.1 Consistency of Initial CPDAG Learning
@@
[…]
@@
B.4.2 Consistency of the Ranking Procedure
@@
Recall from Section 4.2 that the edge-wise independence measure $\widetilde{I}(X,Y)$ for random variables $X,Y$ is calculated by Equation (8) (through discretization). By the proof of Theorem 1, if $\mathcal{G}$ is not a DAG, then there always exists an edge $i-j$ such that $\{X_{i},X_{j}\}$ follows the PANM. Let $\widehat{\widetilde{I}}(X_{i},X_{j})$ be the estimate of $\widetilde{I}(X_{i},X_{j})$ from data. Under the convergence of $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow\widetilde{I}(X_{i},X_{j})$, if random variables $(X_{i},X_{j})$ follow the PANM, then $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow 0$ as $n\rightarrow\infty$. If $(X_{i},X_{j})$ do not follow the PANM, then (A2) and (A3) of Assumption 2 imply that $\widehat{MI}(X_{i},X_{j})\rightarrow MI(X_{i},X_{j})>\delta>0$ and thus $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow\widetilde{I}(X_{i},X_{j})>\frac{\delta}{c_{2}}>0$ as $n\rightarrow\infty$, where $\delta$ is the mutual information lower bound in (A2) and $c_{2}$ is the upper bound of entropy measures in (A3). This reasoning shows that our algorithm would rank all pairs of $(X,Y)$ that satisfy the PANM ahead of those that do not satisfy.
B.4.3 Consistency of Likelihood Ratio Test
@@
Appendix C Additional Results
@@ C.1 Comparison with Nonlinear Constraint-based Algorithms
We compared SNOE with the Randomized Conditional Correlation Test (RCoT) and GAM-based neighborhood selection utilized in CAM. RCoT (Strobl et al., 2019) is a statistical test that approximates the KCI test in faster runtime by utilizing random Fourier features, while the latter regresses all other nodes using a GAM on one node to perform covariate testing and obtain an initial skeleton. We replaced the partial correlation test in the PC algorithm with the RCoT test and GAM-based neighborhood selection separately in SNOE. Effectively, only lines 2 to 2 in Algorithm 2 were changed such that the initial skeleton was learned by these approaches, while the subsequent steps remained the same. A significance level of $\alpha=0.05$ was used in the RCoT test and $\alpha=0.001$ was used in the neighborhood selection approach, which are the default values in their software implementations. We also considered RESIT (Peters et al., 2014), a nonlinear learning method that identifies leaf nodes via kernel-based statistics, but its calculations did not converge. The KCI test (Zhang et al., 2011) and Generalized Covariance Measure, a regression-based CI test proposed in Shah and Peters (2020), were tested too, but failed to converge in a timely manner.
The results are shown in Figure 12. The GAM-regression approach (NbrSel) falls short of the other two, mainly due to a large amount of false positives and false negatives in the skeleton. This demonstrates the necessity of obtaining the correct or a probable parent set for regression, as the resulting model may not accurately approximate the SEM or the covariate selection procedure may be imprecise, and then lead to incorrect inferences. Our approach achieves better performance than RCoT across all function types in all networks, as evidenced by its higher F1 score. It is crucial to note that our algorithm coupled with the RCoT test outperforms NOTEARS, DAGMA, and SCORE, which demonstrates that the test is able to capture nonlinear relations and the advantages of the sequential edge orientation in our algorithm.
@@
C.2 True Positives Captured at Intermediate Steps
The number of true positives captured at each stage of the SNOE algorithm is shown in Figure 13 to complement and explain the intermediate F1 scores. For data generated from nonlinear functions, the edge orientation procedure (step 2) uncovers a great number of true, directed edges from the learned CPDAG. Although the increase in the F1 score is relatively small at this stage since there are additional edges from the candidate set that may be false positives, the F1 score increases greatly once these edges are removed in the third stage. As for the linear, Gaussian DAG, the number of true positives remains unchanged in most cases as the ground truth is the CPDAG, in which its undirected edges should be intact. The results reflect the effectiveness of the edge orientation step in correctly orienting undirected edges. Moreover, the edge pruning step (step 3) deletes few if none true positives.

### v2

@@ B.2 Proof of Proposition 19
$\frac{\frac{1}{n}LR_{n}(\widehat{\theta}_{n},\widehat{\gamma}_{n})}{\widehat{\omega}_{n}/\sqrt{n}}\overset{p}{\longrightarrow}\frac{LR_{n}(\theta^{*},\gamma^{*})}{\sqrt{n}\omega^{*}}\overset{D}{\longrightarrow}N(0,1),\quad\quad n\to\infty,$ (26)
where the first convergence follows from the assumptions on $\log\widehat{F}(x,y\mid z_{1},z_{2})$ and $\log\widehat{G}(x,y\mid z_{1},z_{2})$. Note that $\widehat{\omega}_{n}$ is finite given Assumption A6 in vuong_test. We can characterize its behavior under the large sample limit as $\widehat{\omega}_{n}=O(1)$, since $\widehat{\omega}_{n}\rightarrow\omega^{*},\omega^{*}\in\mathbb{R}$. As $n\rightarrow\infty$, $\frac{1}{n}LR_{n}(\widehat{\theta}_{n},\widehat{\gamma}_{n})$ converges in probability to the true likelihood ratio statistic,
$\frac{1}{n}LR_{n}(\widehat{\theta}_{n},\widehat{\gamma}_{n})\overset{p}{\longrightarrow}\frac{1}{n}LR_{n}({\theta}^{*},{\gamma}^{*})\overset{a.s.}{\longrightarrow}\mathbb{E}\left[\log\frac{F_{\theta^{*}}(X,Y\mid Z_{1},Z_{2})}{G_{\gamma^{*}}(X,Y\mid Z_{1},Z_{2})}\right]\triangleq c.$
When $H_{f}$ is true, $c>0$; when $H_{g}$ is true, $c<0$. All together, the test statistic in Equation (26) can be written as $\sqrt{n}c\cdot O(1)$ and thus converges in probability to $\infty$ or -$\infty$, respectively, under $H_{f}$ or $H_{g}$.
B.3 Proof of Theorem 2
@@
B.3.1 Consistency of Initial CPDAG Learning
@@
Suppose $\{X_{i}\}_{i=1}^{p}$ follows the additive noise model (4) with a DAG $\mathcal{G}_{0}$, which satisfies the conditions for a restricted ANM and (A1) of Assumption 2. Let $\mathcal{E}$ be the CPDAG of $\mathcal{G}_{0}$ and let $\widehat{\mathcal{G}}$ be the graph estimated by the initial learning phase (Lines $\textendash 2$ of Algorithm 2). For some choice of $\alpha_{1},\alpha_{2}\rightarrow 0,P(\widehat{\mathcal{G}}=\mathcal{E})\rightarrow 1$ as $n\rightarrow\infty$.
To prove this lemma, we show that the conditional independence tests are pointwise consistent. In the partial correlation test, let $\rho_{ij\mid S}^{*}$ and $\widehat{\rho}_{ij\mid S}$ respectively denote the true and estimated partial correlation values of $X_{i},X_{j}\mid S]$. Let
$Z_{n}(i,j,S)=\frac{1}{2}\log\frac{1+\widehat{\rho}_{ij\mid S}}{1-\widehat{\rho}_{ij\mid S}}$
be the Fisher z-transformation applied to the estimated partial correlation, and $Z^{*}$ be defined similarly for $\rho_{ij\mid S}^{*}$. Under $H_{0}:X_{i}\perp\!\!\!\!\perp X_{j}\mid S$, the quantity $\sqrt{n-|S|-3}\ Z_{n}(i,j,S)\overset{D}{\longrightarrow}N(0,1)$ as $n\rightarrow\infty$. Let $Z_{1-\gamma_{n}/2}$ denote the critical value corresponding to the percentile $1-\gamma_{n}/2$ under the standard normal distribution. Under this asymptotic distribution, the probability of a type I error $P(|Z_{n}|>Z_{1-\gamma_{n}/2}\mid H_{0})\rightarrow\gamma$ if the sequence $\gamma_{n}\rightarrow\gamma$ as $n\rightarrow\infty$.
@@
B.3.2 Consistency of the Ranking Procedure
@@
Recall from Section 4.2 that the edge-wise independence measure $\widetilde{I}(X,Y)$ for random variables $X,Y$ is calculated by Equation (8) (through discretization). By the proof of Theorem 1, if $\mathcal{G}$ is not a DAG, then there always exists an edge iji-j such that $\{X_{i},X_{j}\}$ follows the PANM. Let $\widehat{\widetilde{I}}(X_{i},X_{j})$ be the estimate of $\widetilde{I}(X_{i},X_{j})$ from data. Under the convergence of $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow\widetilde{I}(X_{i},X_{j})$, if random variables $(X_{i},X_{j})$ follow the PANM, then $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow 0$ as $n\rightarrow\infty$. If $(X_{i},X_{j})$ do not follow the PANM, then (A2) and (A3) of Assumption 2 imply that $\widehat{MI}(X_{i},X_{j})\rightarrow MI(X_{i},X_{j})>\delta>0$ and thus $\widehat{\widetilde{I}}(X_{i},X_{j})\rightarrow\widetilde{I}(X_{i},X_{j})>\frac{\delta}{c_{2}}>0$ as $n\rightarrow\infty$, where $\delta$ is the mutual information lower bound in (A2) and $c_{2}$ is the upper bound of entropy measures in (A3). This reasoning shows that our algorithm would rank all pairs of $(X,Y)$ that satisfy the PANM ahead of those that do not satisfy.
B.3.3 Consistency of Likelihood Ratio Test
@@
Appendix C Supplementary Numerical Results
@@ C.1 Comparison with Nonlinear Constraint-based Algorithms
We compared SNOE with the Randomized Conditional Correlation Test (RCoT) and regression-based neighborhood selection utilized in CAM. RCoT (strobl2019kci_rcot) is a statistical test that approximates the KCI test in faster runtime by utilizing random Fourier features, while the latter method regresses all other nodes using a GAM on one node to perform covariate testing and obtain an initial skeleton. We replaced the partial correlation test in the PC algorithm with the RCoT test and GAM-based neighborhood selection separately in SNOE. Effectively, only lines 2 to 2 in Algorithm 2 were modified such that the initial skeleton was learned by these approaches, while the subsequent steps remained the same. A significance level of $\alpha=0.05$ was used in the RCoT test and $\alpha=0.001$ was used in the neighborhood selection approach, which are the default values in their software implementations. We also implemented the KCI test (kci_test) and generalized covariance measure, a regression-based CI test proposed by shah2020hardness, in the PC algorithm, but both failed to converge in a timely manner. RESIT (resit_anm), a nonlinear learning method that identifies leaf nodes via kernel-based statistics, was considered for comparison but its computations did not converge.
@@
The results are shown in Figure 13. The GAM-regression approach (NbrSel) falls short of the other two, mainly due to a large amount of false positives and false negatives in the skeleton. This demonstrates the benefit of our modified PC algorithm in learning an initial CPDAG. Under the GAM-regression approach, the estimated models may not accurately approximate SEMs or the covariate selection procedure may be imprecise, which led to incorrect inferences of the skeleton and edge orientations. Our approach achieves better performance than RCoT across all function types in all networks, as evidenced by its higher F1 score. It is crucial to note that our algorithm coupled with the RCoT test outperformed NOTEARS, DAGMA, and SCORE (shown in Figure 7), which demonstrates the advantages of the sequential edge orientation in our algorithm paired with methods that can accurately capture nonlinear relations.
C.2 Accuracy of Initial Graph on Algorithm Performance
As SNOE improves upon a learned CPDAG, we investigate how the correctness of the initial graph affects the performance by testing our method separately on the exact, true CPDAG and the learned CPDAG. For the learned CPDAG, we split results by whether the F1 score, calculated with respect to the true CPDAG, is above the median across multiple data sets. The performances of the cross-validation (CV) and sample-splitting (SS) approaches are almost identical, so we only show that of the former. The results are shown in Figure 14, with the F1 score of each approach calculated according to the true DAG for nonlinear functions and the true CPDAG for linear functions.
Figure 14: F1 score of SNOE using different initial graphs: the true CPDAG (‘exact’), a learned CPDAG with above median F1 score (‘AboveMedian’), and a learned CPDAG with below median F1 score (‘BelowMedian’).
We first discuss the linear case when given the exact CPDAG, as one may notice that the F1 score decreases at the final stage. This is because we assume a nonlinear ANM and under this prior assumption, our algorithm extends the PDAG learned in step 3 to a DAG in the final step. Without this assumption, the algorithm would stop after stage 3, where the F1 score generally remains unchanged. Under the other nonlinear settings, we see indeed the initial CPDAG is imperative to learning the true DAG. Yet, the parallel F1 curves between the “AboveMedian” and “BelowMedian” approaches indicate that an inaccurate estimation of a CPDAG does not exacerbate the difference in the final DAG learning accuracy. In fact, the results show that our subsequent steps are useful in learning the true DAG, whether using the true CPDAG or an estimated CPDAG as the starting point.
C.3 True Positives Captured at Intermediate Steps
The number of true positives captured at each stage of the SNOE algorithm is shown in Figure 15 to complement and explain the intermediate F1 scores (see Figure 10). For data generated from nonlinear functions, the edge orientation procedure (step 2) uncovers a great number of true, directed edges from the learned CPDAG. Although the increase in the F1 score is relatively small at this stage since there are additional edges from the candidate set that may be false positives, the F1 score increases greatly once these edges are removed in the third stage. As for the linear, Gaussian DAG, the number of true positives remains unchanged, signaling high precision in the initial graph and correct inference on edge directions in subsequent stages. The results reflect the effectiveness of the edge orientation step in correctly orienting undirected edges. Moreover, the edge pruning step (step 3) deletes few if none true positives.
@@ C.3 True Positives Captured at Intermediate Steps
C.4 Runtime of Intermediate Steps
To breakdown the computational cost, we report the average runtime of each intermediate step in Figure 16. Stages 1 (learning CPDAG by PC algorithm) and 4 (orienting graph into DAG per ANM assumption) exhibit similar runtimes across all networks. We utilized the PC-stable algorithm from the bnlearn package, which implements the iterative testing and orientation process in C++, and chose the partial correlation test for learning the skeleton, hence resulting in a lower runtime. The last stage orients remaining undirected edges in the PDAG, whose runtime is negligible relative to the other three stages. We observe that both stage 2 (edge orientation) and stage 3 (edge deletion) increase with network size, but at an approximately linear rate, indicating that our algorithm is scalable with graph size. SNOE-CV has a slightly higher runtime in stage 2 since we employ cross validation in the likelihood test. Moreover, the use of Meek’s rules assists in further detecting more directed edges, hence reducing the actual number of edges to evaluate.
Figure 16: Average runtime of intermediate steps by network size. The steps are (1) initial CPDAG learning, (2) edge orientation, (3) edge deletion, and (4) graph refinement.
C.5 Sensitivity to the Misranking of Edges
Given that the learned CPDAG is usually not perfect, it is difficult to supply a correct ranking of the undirected edges to the algorithm. Instead, we compare the results of orienting edges by our ranking as in SNOE-SS and SNOE-CV to those of orienting edges in an arbitrary order (SNOE-SS-Rand, SNOE-CV-Rand). We apply the approaches to data sets generated under linear, invertible(inv), and non-invertible(ninv) functions with Gaussian noise. The simulated data sets use the exact settings described in Section 6.3. Then, we examine F1 scores after the orientation stage of two select networks (mehra and magic2) in Figure 7 under the different functions. Table 4 shows the F1 score after orienting edges in a random or ranked order, i.e. after stage 2 in Figure 10.
network approach F1_linear F1_inv F1_ninv
mehra SNOE-CV 0.516 0.551 0.504
mehra SNOE-CV-Rand 0.514 0.536 0.504
mehra SNOE-$S$ 0.524 0.554 0.510
mehra SNOE-$S$-Rand 0.512 0.533 0.499
magic2 SNOE-CV 0.761 0.631 0.639
magic2 SNOE-CV-Rand 0.744 0.628 0.620
magic2 SNOE-$S$ 0.766 0.632 0.631
magic2 SNOE-$S$-Rand 0.753 0.618 0.612
Table 4: F1 scores of the graphs after evaluating edges in a ranked or random manner in the orientation stage.
All approaches are applied to the same initial graph. The edge ranking procedure indeed led to more true positives in the orientation stage, as evidenced by the higher F1 scores. One may notice that the difference between the ranked and random procedures is not large. Under a random order, a greater proportion of undirected edges not satisfying the PANM criterion, due to missing parent variables, are evaluated by the likelihood test. As the algorithm cannot accurately estimate the true regression function due to missing parent variables, the likelihood test would likely find models for $X\rightarrow Y$ and $Y\rightarrow X$ equivalent, and thus keep the edge undirected. These edges will be correctly oriented later, once they meet the PANM condition following the orientation of other edges incident on the two nodes. In summary, the ranking procedure offers an improvement over a random ordering of edges. In the case where edges are misranked, the likelihood test prevents incorrect orientation and mitigates the impact of misranking.